solve:
2y(y+4)=-5(y+3)
2y(y+4)-5(y+3)=0
unclear as to where to take it from there.
I don't see how they got -3/2,5 out of that. I'm trying though.

Answer by uuuu(3)

use the quadratic formula for the problem 2y^2+3y-15=0
I might be a little off since im doing this in my head but you should get close to or exactly y=-3/4-1/4 square root of 129 and y=-3/4+1/4 square root of 129 if you need to see the steps if youre still not getting it email me back